# Aperiodic Structures Never Collapse: Fibonacci Hierarchies for Lossless Compression

### 16 Mar 2026 v1

@@ Abstract
We validate these results with Quasicryth, [^1] Quasicryth: portmanteau of quasicrystal and compression; the name also reflects the internal structure — quasicrystalline tiling hierarchy. a lossless text compressor built on a ten-level Fibonacci hierarchy with phrase lengths $\{2,3,5,8,13,21,34,55,89,144\}$ words. In controlled A/B experiments with identical codebooks, the aperiodic advantage over Period-5 grows from 1,372 B at 3 MB to 1,349,371 B at 1 GB, explained by the activation of deeper hierarchy levels. On enwik9 (1 GB), Quasicryth achieves 359,883,431 B (35.99%), with 45,608,715 B attributable to the quasicrystal tiling alone.
@@ 1. Introduction
    Deep hierarchy with ten levels. Quasicryth v5.5 implements the full Fibonacci substitution hierarchy to depth $k=9$, providing lookup positions for phrases of up to 144 words. At enwik9 scale (298 M words), the 89-gram and 144-gram levels activate for the first time, contributing 2,396 and 945 deep hits respectively — positions entirely unavailable to any periodic tiling.
@@
    Quantified aperiodic advantage at scale. Controlled A/B experiments: at 100 MB the Fibonacci tiling saves 40,385 B over Period-5; at 1 GB this grows to 1,349,371 B. The 89-gram and 144-gram levels activate at 1 GB (enwik9), contributing 2,396 and 945 deep hits — positions structurally unavailable to any periodic tiling.
@@
    Single-parameter tiling. The entire parsing structure — all tile boundaries, hierarchy levels, deep n-gram positions, and adaptive model assignments — is determined by a single 2-byte phase parameter. The decoder regenerates the complete tiling deterministically; no structural information appears in the compressed bitstream.
@@
    Separate bz2 escape stream. Out-of-vocabulary words are segregated into a parallel stream compressed with bzip2 [22], achieving $\sim$2.6 bpb versus $\sim$3.5 bpb for inline arithmetic coding. The quasicrystalline tiling governs which words escape, maintaining the structural integrity of both streams.
@@ 3. Method
Quasicryth processes input text through an eight-step pipeline. All design choices reflect the constraint that the tiling structure must be perfectly recoverable from a single phase parameter.
@@ 3.1 Step 1: Case Separation
The input byte stream is tokenised into words and punctuation tokens. Each token is classified as lowercase, titlecase, or uppercase, yielding a 3-symbol case flag stream encoded separately with a 18-bit-precision arithmetic coder (256-symbol adaptive model with periodic rescaling). The remaining pipeline operates on the fully lowercased token stream. On enwik9, case coding contributes 21,315,698 B ($\approx$2.13% of original), a necessary overhead for exact roundtrip.
@@ 3.3 Step 3: Multi-Level Codebook Construction
Eleven frequency-ranked codebooks are built from the word token sequence, corresponding to phrase lengths $\{1,2,3,5,8,13,21,34,55,89,144\}$ — the first eleven Fibonacci numbers. Table 1 shows the codebook sizes. All n-gram codebooks are filtered to entries where every constituent word appears in the unigram codebook. For n-grams with $n\geq 13$, frequency counters are periodically pruned (singletons removed every 500 K entries) to prevent out-of-memory on large inputs. The serialised codebook is zlib-compressed; at enwik9 scale this occupies only 77,585 B.
@@
Unigram 1 4,000 8,000 8,000
Bigram 2 3,000 6,000 6,000
Trigram 3 1,000 2,000 2,000
5-gram 5 1,000 2,000 2,000
8-gram 8 500 1,000 1,000
13-gram 13 300 1,000 1,000
21-gram 21 100 500 500
34-gram 34 50 200 500
55-gram 55 0 0 200
89-gram 89 0 0 200
144-gram 144 0 0 100
Index encoding within the arithmetic coding stream uses a two-stage scheme: indices 0253253 are encoded directly (1 AC symbol); index $254$ signals an extended lookup (next symbol gives 254254508508); index $255$ signals an out-of-vocabulary escape.
@@ 3.4 Step 4: Quasicrystalline Word-Level Tiling
The one-dimensional Fibonacci quasicrystal is generated via the cut-and-project method. For tile position $k$ and phase $\theta\in[0,1)$:
$\text{tile}(k)=\begin{cases}L&\text{if }\lfloor(k+1+\theta)\,/\,\varphi\rfloor-\lfloor(k+\theta)\,/\,\varphi\rfloor=1\\ S&\text{otherwise}\end{cases}$ (1)
where $\varphi=(1+\sqrt{5})/2$ is the golden ratio. Each $L$ tile consumes 2 words (bigram lookup); each $S$ tile consumes 1 word (unigram lookup). The quasicrystalline matching rule — no two $S$ tiles adjacent — is enforced by merging any resulting $SS$ pair into an $L$ tile.
The compressor evaluates 32 equally spaced phases $\theta\in\{0/32,1/32,\ldots,31/32\}$, selecting the phase that maximises the scoring function:
@@ 3.5 Step 5: Deep Substitution Hierarchy Construction
At each deflation level $k$, the super-L tile at that level spans $F_{k+2}$ words, where $F_{n}$ is the $n$-th Fibonacci number. The function detect_deep_positions() traces each $L$ tile upward through the hierarchy: a tile at position $t_{i}$ can attempt a level-$k$ n-gram lookup if and only if it is the leftmost child of a super-L at every level up to $k$, and the span covers exactly $F_{k+2}$ words.
@@ 3.6 Step 6: Multi-Level Adaptive Arithmetic Coding
Encoding proceeds tile-by-tile. At each $L$ tile position, the deepest applicable hierarchy level is tried first. For level $k$ from $k_{\max}$ down to $1$: if the tile is at a super-L position at level $k$, encode a binary flag (hit/miss) via model lvl_flag[$k$]. On a hit, encode the codebook index via models lvl_idx[$k$] and lvl_ext[$k$], and advance the tile pointer by $F_{k+2}$ words. On a miss, fall through to the next lower level. After all deep levels fail, encode the bigram index via L_model[$c$] where $c\in\{0,\ldots,7\}$ is a 3-bit hierarchy context derived deterministically from the tile’s position in its parent, parent type, and grandparent type. $S$ tiles are encoded analogously via S_model[$c$].
Algorithm 1 summarises the encoding procedure. All arithmetic coding models use 256-symbol adaptive tables with 24-bit register precision and periodic frequency rescaling (halve all counts when the total exceeds 16,384).
@@
0: Tile array $T$, word array $W$, codebooks $\mathcal{C}$
1: Build substitution hierarchy from $T$
2: Detect deep positions: $\text{can}[k][t_{i}]$ for $k=1\ldots k_{\max}$
@@
4:while $i<|T|$ do
5:  if $T[i]=L$ then
6:   for $k=k_{\max}$ downto 11 do
7:    if $\text{can}[k][i]$ then
8:     Look up $W[i..i{+}F_{k+2}{-}1]$ in $\mathcal{C}_{k}$
9:     Encode flag $\mathbf{1}$ via lvl_flag[$k$]
10:     if hit then
11:      Encode index; $i\leftarrow i+F_{k+2}$; goto next
12:     end if
13:     Encode flag $\mathbf{0}$
14:    end if
15:   end for
16:   Encode bigram (W[i],W[i+1])(W[i],W[i{+}1]) via L_model[$c(i)$]
17:   $i\leftarrow i+2$
18:  else
19:   Encode unigram $W[i]$ via S_model[$c(i)$]
20:   $i\leftarrow i+1$
21:  end if
22:end while
3.7 Step 7: Separate bz2 Escape Stream
Out-of-vocabulary words (codebook misses) are collected into a separate buffer as length-prefixed raw byte sequences. The buffer is compressed with bzip2 (BWT + Huffman coding). This dual-stream architecture is critical: the arithmetic coding stream contains only compact integer indices and binary flags, while rare words compress efficiently under BWT due to their natural lexicographic clustering. On enwik9, the AC payload is 254,952,735 B and the bz2 escape stream is 83,537,358 B — combined they account for 93.6% of the total compressed size.
@@ 3.8 Step 8: Output Assembly
Codebook (zlib) variable
Escape stream (bz2) variable
@@
The complete parsing structure — tile boundaries, hierarchy depths, deep n-gram positions, and model assignments — is implicit in the 2-byte phase value. The decompressor regenerates the Fibonacci tiling from $\theta$, rebuilds the substitution hierarchy, and decodes the AC stream identically, with no structural side-channel required.
@@ 4.4 Theorem: Fibonacci Hierarchy Never Collapses
The frequency vector of the Fibonacci tiling is exactly the Perron-Frobenius eigenvector $\mathbf{v}_{1}$ (Eq. (7)). Under $(M^{-1})^{k}$:
@@
The frequency vector is scaled by $1/\varphi^{k}$ but its direction is preserved exactly. The ratio $n_{L}^{(k)}/n_{S}^{(k)}=\varphi>0$ persists at every level $k$. Since $n_{L}^{(k)},n_{S}^{(k)}>0$ for all finite $k$, both supertile types are always present. ∎
@@ 4.5 Theorem: All Periodic Tilings Collapse
Decompose the frequency vector in the eigenbasis of $M$:
@@
where $\mathbf{v}_{2}$ is the eigenvector for $\lambda_{2}=\psi$. Since $\mathbf{v}$ has rational entries and $\mathbf{v}_{1}$ has irrational entries (containing $\sqrt{5}$), we must have $\beta\neq 0$. Under $(M^{-1})^{k}$:
@@
Since $|\psi^{-1}|=\varphi>1$, the second component grows as $\varphi^{k}$, dominating for large $k$. The frequency trajectory $n_{L}^{(k)}/n_{S}^{(k)}$ therefore diverges from $\varphi$: it oscillates and eventually sends one count to zero (or negative, indicating collapse has already occurred). The collapse timescale is determined by when $\beta\varphi^{k}|\psi^{-k}|$ exceeds the first component:
$k^{*}\approx\frac{\log p}{\log\varphi}$ (15)
since the deviation $|\beta|\propto 1/p$ for the best periodic approximant of $\varphi$ with period $p$. ∎

### 23 Mar 2026 v2

@@ Abstract
We validate these results with Quasicryth, [^1] Quasicryth: portmanteau of quasicrystal and compression; the name also reflects the internal structure — quasicrystalline tiling hierarchy. Source code: https://github.com/robtacconelli/quasicryth. a lossless text compressor built on a ten-level Fibonacci hierarchy with phrase lengths $\{2,3,5,8,13,21,34,55,89,144\}$ words. Version 5.6 extends the tiling engine with 36 multi-structure tilings (12 golden-ratio Fibonacci phases plus 6 original non-golden irrational tilings and 18 optimized additions discovered by iterative greedy alpha search, including the far-out $\alpha=0.502$), scanning internally to select the best tiling per block; no tiling parameter appears in the compressed header. In controlled A/B experiments with identical codebooks, the aperiodic advantage over Period-5 grows from 36,243 B at 3 MB to 11,089,469 B at 1 GB, explained by the activation of deeper hierarchy levels. On enwik9 (1 GB), Quasicryth achieves 225,918,349 B (22.59%) in multi-structure mode, with 20,735,733 B saved by the Fibonacci tiling relative to no tiling.
@@ 1. Introduction
    Deep hierarchy with ten levels. Quasicryth v5.6 implements the full Fibonacci substitution hierarchy to depth $k=9$, providing lookup positions for phrases of up to 144 words. At enwik9 scale (298 M words), the 89-gram and 144-gram levels activate for the first time, contributing 5,369 and 2,026 deep hits respectively — positions entirely unavailable to any periodic tiling.
@@
    Quantified aperiodic advantage at scale. Controlled A/B experiments: at 3 MB the Fibonacci tiling saves 36,243 B over Period-5; at 1 GB this grows to 11,089,469 B. The 89-gram and 144-gram levels activate at 1 GB (enwik9), contributing 5,369 and 2,026 deep hits — positions structurally unavailable to any periodic tiling.
@@
    Multi-structure tiling engine. Version 5.6 extends the Fibonacci-only design to 36 tilings: 12 golden-ratio Fibonacci phases, 6 original non-golden irrational tilings, and 18 optimized additions discovered by iterative greedy alpha search (including $\alpha=0.502$, far below the golden ratio, which contributes massive trigram/5-gram coverage). The encoder scans all 36 candidates internally and selects the best tiling per block; the decoder regenerates the complete tiling deterministically without any tiling parameter in the compressed header.
@@
    Competitive with established compressors. Quasicryth now surpasses bzip2 and approaches xz-level compression on large corpora, demonstrating that quasicrystalline tiling hierarchies are a practical compression engine, not merely a theoretical curiosity.
  8. 8.
    Separate LZMA escape stream. Out-of-vocabulary words are segregated into a parallel stream compressed with LZMA, achieving strong byte-level compression versus inline arithmetic coding. The quasicrystalline tiling governs which words escape, maintaining the structural integrity of both streams.
@@ 3. Method
Quasicryth processes input text through an eight-step pipeline.
@@ 3.1 Step 1: Case Separation
The input byte stream is tokenised into words and punctuation tokens. Each token is classified as lowercase, titlecase, or uppercase, yielding a 3-symbol case flag stream encoded separately with a 24-bit-precision adaptive arithmetic coder with order-2 context (9 contexts from previous two case flags, variable-alphabet Fenwick-tree models). The remaining pipeline operates on the fully lowercased token stream. On enwik9, case coding contributes 20,397,073 B ($\approx$2.04% of original), a necessary overhead for exact roundtrip.
@@ 3.3 Step 3: Multi-Level Codebook Construction
Eleven frequency-ranked codebooks are built from the word token sequence, corresponding to phrase lengths $\{1,2,3,5,8,13,21,34,55,89,144\}$ — the first eleven Fibonacci numbers. Table 1 shows the codebook sizes. All n-gram codebooks are filtered to entries where every constituent word appears in the unigram codebook. For n-grams with $n\geq 13$, frequency counters are periodically pruned (singletons removed every 500 K entries) to prevent out-of-memory on large inputs. The serialised codebook is LZMA-compressed; at enwik9 scale this occupies only 533,680 B.
@@
Unigram 1 8,000 16,000 64,000
Bigram 2 4,000 8,000 32,000
Trigram 3 2,000 4,000 32,000
5-gram 5 1,000 2,000 16,000
8-gram 8 500 1,000 4,000
13-gram 13 300 1,000 4,000
21-gram 21 100 500 2,000
34-gram 34 50 200 2,000
55-gram 55 0 100 1,000
89-gram 89 0 0 1,000
144-gram 144 0 0 500
Index encoding uses variable-alphabet adaptive arithmetic coding with Fenwick-tree acceleration, supporting codebook sizes up to 64,000 entries per level.
@@ 3.4 Step 4: Quasicrystalline Word-Level Tiling
v5.6 employs 36 aperiodic tilings from multiple families of irrational numbers: 12 golden-ratio ($\alpha=1/\varphi$) tilings with phases equally spaced by $1/\varphi$, plus 6 original non-golden tilings (2 each from $\alpha=\sqrt{58}-7$, noble-5, and $\sqrt{13}-3$), and 18 optimized additions discovered by iterative greedy alpha optimization. The greedy search evaluates candidate $\alpha$ values by measuring the marginal deep-position gain over existing tilings, then commits the best candidate and repeats. This process discovered far-out irrationals such as $\alpha=0.502$ (well below the golden ratio $1/\varphi\approx 0.618$), which provides massive trigram and 5-gram coverage at positions complementary to all golden-ratio tilings. All 36 tilings produce quasicrystalline L/S sequences via the cut-and-project method.
For tile position $k$, irrational slope $\alpha$, and phase $\theta\in[0,1)$:
$\text{tile}(k)=\begin{cases}L&\text{if }\lfloor(k+1)\alpha+\theta\rfloor-\lfloor k\alpha+\theta\rfloor=1\\ S&\text{otherwise}\end{cases}$ (1)
When $\alpha=1/\varphi$ (golden ratio), this reduces to the classical Fibonacci quasicrystal. Each $L$ tile consumes 2 words (bigram lookup); each $S$ tile consumes 1 word (unigram lookup). The quasicrystalline matching rule — no two $S$ tiles adjacent — is enforced by merging any resulting $SS$ pair into an $L$ tile.
A command-line flag -f restricts to the 12 golden-ratio tilings only (Fibonacci-only mode) for controlled A/B comparison.
The compressor evaluates all 36 tilings, selecting the tiling that maximises the scoring function:
@@ 3.5 Step 5: Deep Substitution Hierarchy Construction
At each deflation level $k$, the super-L tile at that level spans $F_{k+2}$ words, where $F_{n}$ is the $n$-th Fibonacci number. The function detect_deep_positions() traces each $L$ tile upward through the hierarchy: a tile at position $t_{i}$ can attempt a level-$k$ n-gram lookup if and only if it is the leftmost child of a super-L at every level up to $k$, and the span covers exactly $F_{k+2}$ words. Deep matches from all 36 tilings are collected into position-indexed arrays, then a greedy non-overlapping selection (deepest match wins) assigns each word position to its optimal encoding level.
@@ 3.6 Step 6: Multi-Level Adaptive Arithmetic Coding
Encoding proceeds event-by-event over the greedy-selected assignment. Five interacting mechanisms produce the final bitstream:
Order-2 level context.
Level decisions use an order-2 context model conditioned on $(\mathit{prev\_level},\mathit{prev\_prev\_level})$, yielding $12\times 12=144$ specialised 12-symbol adaptive models.
Context-conditioned indices.
Index models are conditioned on the previous level, giving 12 variants per codebook level.
Recency cache.
A 64-entry recency cache per encoding level implements move-to-front caching. Each index event first checks the cache; on a hit, only the cache position is encoded (much cheaper than the full index).
Two-tier unigram.
[…]
Word-level LZ77.
A word-level LZ77 pass scans the event stream for repeated sequences of 3+ consecutive (level, index) tuples. Matches are encoded as (offset, length) pairs with log-scale offset coding, replacing the individual AC symbols for matched events.
Algorithm 1 summarises the encoding procedure. All arithmetic coding models use variable-alphabet Fenwick-tree adaptive tables with 24-bit register precision and periodic frequency rescaling (halve all counts when the total exceeds 16,384).
@@
0: Event array $E$ (from greedy selection), word array $W$, codebooks $\mathcal{C}$
1: Run word-level LZ77 over $E$; mark matched spans
2: Initialise 144 level-context models, 12 per-level index models, caches
@@
4:for each event $e_{i}$ in $E$ do
5:  if $e_{i}$ is LZ77 match start then
6:   Encode LZ77 flag, offset, length
7:   Skip matched events; update $\mathit{pl}_{1},\mathit{pl}_{2}$
8:  else
9:   Encode level $\ell_{i}$ via level model[$\mathit{pl}_{2}$$\mathit{pl}_{1}$]
10:   if $\ell_{i}=0$ (unigram) then
11:    Check cache[$\ell_{i}$]; if hit, encode cache position
12:    Else: encode tier flag, then index via tier model
13:   else
14:    Check cache[$\ell_{i}$]; if hit, encode cache position
15:    Else: encode index via idx_model[$\ell_{i}$$\mathit{pl}_{1}$]
16:   end if
17:   Update cache[$\ell_{i}$] with move-to-front
18:   $\mathit{pl}_{2}\leftarrow\mathit{pl}_{1}$; $\mathit{pl}_{1}\leftarrow\ell_{i}$
19:  end if
20:end for
3.7 Step 7: Separate LZMA Escape Stream
Out-of-vocabulary words (codebook misses) are collected into a separate buffer as length-prefixed raw byte sequences. The buffer is compressed with LZMA. This dual-stream architecture is critical: the arithmetic coding stream contains only compact integer indices and binary flags, while rare words compress efficiently under LZMA due to their natural lexicographic clustering. On enwik9, the AC payload is 180,760,315 B and the LZMA escape stream is 24,227,220 B — combined they account for 90.7% of the total compressed size.
@@ 3.8 Step 8: Output Assembly
Codebook (LZMA) variable
Escape stream (LZMA) variable
@@
The decompressor reads the header fields, reconstructs the tiling and hierarchy, and decodes the AC stream identically, with no structural side-channel required.
@@ 4.4 Theorem: Fibonacci Hierarchy Never Collapses
The frequency vector of the Fibonacci tiling is exactly the Perron-Frobenius eigenvector $\mathbf{v}_{1}$ (Eq. (7)). Since $\mathbf{v}_{1}$ is an eigenvector of $M$ with eigenvalue $\varphi$, it is also an eigenvector of $M^{-1}$ with eigenvalue $1/\varphi$. Applying the deflation matrix $k$ times:
@@
Explicit computation at the first five deflation levels.  Start with a Fibonacci tiling of $N$ tiles. The initial count vector is $\mathbf{n}^{(0)}=N\,\mathbf{v}_{1}=\frac{N}{\varphi+1}\bigl(\varphi,\;1\bigr)^{T}$. Applying the deflation rule $n_{L}^{(k+1)}=n_{S}^{(k)}$, $n_{S}^{(k+1)}=n_{L}^{(k)}-n_{S}^{(k)}$ (Eqs. (10)–(11)), the count vectors at levels $k=0,\ldots,4$ are:
$k$ $(n_{L},\,n_{S})$ normalised $n_{L}/n_{S}$ Scale
0 (0.618, 0.382)(0.618,\;0.382) $\varphi$ 11
1 (0.382, 0.236)(0.382,\;0.236) $\varphi$ $0.618$
2 (0.236, 0.146)(0.236,\;0.146) $\varphi$ $0.382$
3 (0.146, 0.090)(0.146,\;0.090) $\varphi$ $0.236$
4 (0.090, 0.056)(0.090,\;0.056) $\varphi$ $0.146$
At every level the count vector remains proportional to $(\varphi,\;1)$. The total number of supertiles shrinks by a factor of $1/\varphi$ per level (column 4), but the ratio $n_{L}/n_{S}=\varphi$ is preserved exactly — not approximately. This is because $\mathbf{v}_{1}$ is an eigenvector of $M^{-1}$: multiplying an eigenvector by its matrix simply rescales it, without rotating its direction.
Concretely, for enwik9 (W=298,263,298W=298{,}263{,}298 words, $N\approx 114{,}000{,}000$ tiles), the count vectors are approximately:
$k$ $n_{L}^{(k)}$ $n_{S}^{(k)}$ $n_{L}/n_{S}$
0 70,451,00070{,}451{,}000 43,549,00043{,}549{,}000 $1.61803\ldots$
1 43,549,00043{,}549{,}000 26,902,00026{,}902{,}000 $1.61803\ldots$
2 26,902,00026{,}902{,}000 16,626,00016{,}626{,}000 $1.61803\ldots$
3 16,626,00016{,}626{,}000 10,276,00010{,}276{,}000 $1.61803\ldots$
4 10,276,00010{,}276{,}000 6,350,0006{,}350{,}000 $1.61803\ldots$
At level 4, there are still over 16 million supertiles, with both types abundantly present. The ratio remains exactly $\varphi$ at every level.
Why both counts remain positive.  Since the ratio $n_{L}^{(k)}/n_{S}^{(k)}=\varphi>0$ is a fixed positive constant for all $k$, and the total count $n_{L}^{(k)}+n_{S}^{(k)}>0$ for any finite tiling, it follows that both $n_{L}^{(k)}>0$ and $n_{S}^{(k)}>0$ individually. Neither count can reach zero because that would require the ratio to become either $\infty$ or 0, which contradicts $n_{L}^{(k)}/n_{S}^{(k)}=\varphi$ for all $k$. ∎
@@ 4.5 Theorem: All Periodic Tilings Collapse
The eigenvectors of $M=\bigl(\begin{smallmatrix}1&1\\ 1&0\end{smallmatrix}\bigr)$ for eigenvalues $\lambda_{1}=\varphi$ and $\lambda_{2}=\psi$ are:
$\mathbf{v}_{1}=\frac{1}{\varphi+1}\begin{pmatrix}\varphi\\ 1\end{pmatrix},\quad\mathbf{v}_{2}=\frac{1}{\psi+1}\begin{pmatrix}\psi\\ 1\end{pmatrix}$,
where $\psi+1=(3{-}\sqrt{5})/2\approx 0.382$.
Step 1: Decompose the frequency vector in the eigenbasis.  Any frequency vector $\mathbf{v}=(r,\;1-r)^{T}$ with $r=n_{L}/p$ can be written as
@@
where $\mathbf{v}_{2}$ is the eigenvector for $\lambda_{2}=\psi$. Since $\mathbf{v}$ has rational entries and $\mathbf{v}_{1}$ has irrational entries (containing $\sqrt{5}$), we must have $\beta\neq 0$.
To find $\alpha$ and $\beta$, we solve the $2\times 2$ system. For the second components: $1-r=\alpha/(\varphi+1)+\beta/(\psi+1)$. For the first components: $r=\alpha\varphi/(\varphi+1)+\beta\psi/(\psi+1)$. The solution is:
$\alpha=\frac{r-\psi(1-r)}{(\varphi-\psi)/(\varphi+1)}=\frac{(\varphi+1)\bigl[r-\psi(1-r)\bigr]}{\sqrt{5}},$
and similarly for $\beta$ with $\varphi$ and $\psi$ interchanged. The key point is that $\beta=0$ if and only if $r/(1-r)=\varphi$, i.e., $r=\varphi/(\varphi+1)$, which is irrational. So for any rational $r$ we have $\beta\neq 0$.
Step 2: Explicit computation for Period-5.  Period-5 has $r=3/5$, so $\mathbf{v}=(3/5,\;2/5)^{T}$. We solve $(3/5,\;2/5)^{T}=\alpha\,\mathbf{v}_{1}+\beta\,\mathbf{v}_{2}$. Using $\varphi=1.61803\ldots$, $\psi=-0.61803\ldots$, $\varphi+1=2.61803\ldots$, $\psi+1=0.38197\ldots$:
$\alpha$ $=\frac{(\varphi{+}1)[3/5-\psi\cdot 2/5]}{\sqrt{5}}$
$=\frac{2.618\times 0.847}{2.236}\approx 0.992,$
$\beta$ $=\frac{(\psi{+}1)[\varphi\cdot 2/5-3/5]}{\sqrt{5}}$
$=\frac{0.382\times 0.047}{2.236}\approx 0.008.$
The coefficient $\beta$ is small (Period-5 is a good approximant of $\varphi$) but crucially nonzero.
Step 3: Apply $(M^{-1})^{k}$ and watch the divergence.  Under $(M^{-1})^{k}$:
@@
Since $|\psi^{-1}|=\varphi>1$, the second component grows as $\varphi^{k}$, dominating for large $k$. Let us track both components for $k=0,1,2,3,4$:
$k$ $\alpha/\varphi^{k}$ $\beta/\psi^{k}$ $(n_{L},n_{S})$ $n_{L}/n_{S}$
0 0.992 0.008 $(3,2)$ 1.50
1 0.613 -0.013 $(2,1)$ 2.00
2 0.379 0.021 $(1,2)$ 0.50
3 0.234 -0.034 $(1,0)$ $\infty$
4 0.145 0.055 $(0,1)$ 0
At level 3, all tiles are L (ratio $\infty$); at level 4, all tiles are S (ratio =0=0). In both cases one tile type has vanished — collapse has occurred. The $\beta$ component, though initially tiny ($0.008$), grows by $|\psi^{-1}|=\varphi\approx 1.618$ per level, while the $\alpha$ component shrinks by $1/\varphi$ per level. By level 3, the growing component has overwhelmed the shrinking one.
Step 4: General collapse timescale.  The frequency trajectory $n_{L}^{(k)}/n_{S}^{(k)}$ therefore diverges from $\varphi$: it oscillates and eventually sends one count to zero (or negative, indicating collapse has already occurred). The collapse timescale is determined by when $|\beta|\,\varphi^{k}$ exceeds $\alpha/\varphi^{k}$, i.e. when $|\beta/\alpha|\,\varphi^{2k}\geq 1$:
$k^{*}\approx\frac{\log(\alpha/|\beta|)}{2\log\varphi}\approx\frac{\log p}{\log\varphi}$ (15)
since the deviation $|\beta|\propto 1/p$ for the best periodic approximant of $\varphi$ with period $p$. For Period-5: $\log 5/\log\varphi\approx 3.3$, consistent with collapse first occurring at level 3 (all-L) and completing at level 4 (all-S). ∎